900 people attended a dance. The cost of admission was $2.00 per person or $3.50 per couple. Twice as many $2.00 tickets as $3.50 were sold. Find the amount of money obtained from ticket sales?
Found 2 solutions by ewatrrr, edjones:
Answer by ewatrrr(24785)


Hi,
Let x represent the amount of couple tickets, then 2x would be the amount of single tickets
2x + x(two people per ticket) = 2x + 2x = 900 people
4x =900
x = 225, the amount of couple tickets, 450 the number of single tickets
Money from tickets sales
450*$2.00 + 225*$3.50 =

Answer by edjones(8007)

900 Let s=number of singles and c=number of couples
Twice as many singles as couples were there.
450 people were singles and bought 450 singles tickets.
450 people were couples and bought 225 couples tickets.
450+225=675 tickets sold.
450*2.00=$900
225*3.50=$787.50
-----------------add
========>$1687.50 the amount of money obtained from ticket sales.